AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Thu Apr 3 19:29:11 2025
Author
*
:
Subject
*
:
<p><strong>14:45</strong></p> <p>New run 55 with:</p> <p>Right arm at 64.60 mm</p> <p>Left arm IN at 36.54 mm</p> <p>MBS working - hopefully</p> <p>Target 7E11</p> <p>Injection ca. 5E7</p> <p>horizontal beam position 8.75 mm</p> <p>vertical beam position -5mm</p> <p>Changed electron cooler set point to 244.2 V</p> <p>Read value 243.66 V</p> <p>Collector current 4.31A</p> <p>Rates OK. Temps OK.</p> <p> </p> <p><strong>15:00</strong></p> <p>Detector 0 was noisy. Shook arm. Fixed it.</p> <p> </p> <p><strong>16:15</strong></p> <p>Detector 0 was noisy. Shook arm. Fixed it.</p> <p>There is no beam in CRYRING since 16:05.</p> <p> </p> <p><strong>16:25</strong></p> <p>The beam is back but it is still a bit unstable and the operators are still trying optimize it.</p> <p> </p> <p><strong>17:20</strong></p> <p>We just did standard detector/electronics checks. Everthing looks OK.</p> <p> </p> <p><strong>17:50</strong></p> <p>Take screenshots.</p> <p> </p> <p><strong>20:50</strong></p> <p>Detector 0 was noisy. Shook arm. Fixed it.</p> <p>But now detector 2 seems to have lost the Rutherford peak.</p> <p>Screenshots taken and attached.</p> <p> </p> <p><strong>21:05</strong></p> <p>We did "Check ASIC Control" following Carlo's advise. This fixed the detector 2, but the detector 0 got noisy again. Another movement of the detector fixed the noise again.</p> <p> </p> <p><strong>22:15</strong></p> <p>Detector 0 was noisy. Shook arm. Fixed it.</p> <p> </p> <p><strong>23:05</strong></p> <p>This run has finished. The cooler voltage will be increased by 0.1V for the next run.</p> <p>The latest spectra and detector/DAQ status are in the attachment.</p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown